# Towards Invisible Backdoor Attack on Text-to-Image Diffusion Model

### v1

@@ 4 Experiments
IBA (Ours) 97.50 3.00 0.50 0.75 20.82
Table 1: Comparison of the proposed method with current methods on Attack Success Rate (ASR), Detect Success Rate (DSR) and Frechet Inception Distance (FID). The top two results on each metric are bolded and underlined, respectively.
@@
4.1 Settings
Attack configurations and baselines. We select the syntactic template (DET) (NOUN) (ADP) (DET) (NOUN) (VERB) (ADP) (NOUN) as the trigger for the following experiment, as it has the lowest frequency on DiffusionDB [45] while still producing human-readable backdoor prompts. We leverage GPT-4o [3] to generate a total of 1100 backdoor samples for each trigger. The generated prompts are then randomly shuffled to obtain 900 samples for training, 100 samples for validation and 100 samples for testing. During training, we use the AdamW [26] optimizer with a learning rate of 1e-4 and train the victim model for 600 epochs. The hyperparameters $\gamma$ and $\lambda$ are set to 1 and 0.01, respectively. We compare with four state-of-the-art methods on backdoor attack, including Rickrolling [41], Villan Diffusion [8], EvilEdit [43] and BadT2I [48]. Training details of each baseline are provided in the Sec. A.3. Following the original settings [41, 44], we use the stable diffusion v1.4 (sd14) [34] as the victim model.
Evaluation Settings. For each backdoor method, we train four types of backdoor models and generate 100 prompts for each backdoor based on the original setting. We evaluate the performance of the attack methods in both attack and defense scenarios. For the attack scenario, we compute two metrics: 1) Frechet Inception Distance (FID) [30], which reflects the model’s ability to maintain performance on benign samples. We compute the FID using the COCO-30k validation subset [23]. 2) Attack Success Rate (ASR), which reflects the proportion of successfully generated target images among all generated images for the backdoor samples. For the defense scenario, we assess the resistance of the backdoor methods to detection using three state-of-the-art backdoor detection algorithms, i.e. T2ISheild-FTT [44], T2ISheild-CDA [44] and UFID [15]. We measure the Detection Success Rate (DSR) of these algorithms on the test samples to evaluate the resistance of the backdoor attack methods to detection. The detailed parameters for each defense methods are listed in Sec. A.3.
4.2 Qualitative Results
@@
Fig. 3 shows the attention visualization results for benign samples and backdoor samples from models trained with the Rickrolling [41] and our IBA. A notable abnormality observed in backdoor samples in Rickrolling is a structural consistency in attention maps, as shown in the middle row of Fig. 3. In contrast, the model trained with IBA reduces this consistency while still effectively generating attacker-specified content. In Fig. 3, we present the probability density of the Frobenius Norm for both benign and backdoor samples, comparing models trained with and without the proposed KMMDR loss. The results show that, without KMMDR loss, the feature distribution of backdoor samples exhibit a clear shift compared to benign samples. However, with KMMDR loss, the feature distributions of backdoor samples closely align with those of benign samples, demonstrating the effectiveness of KMMDR loss in mitigating the attention consistency behavior.
@@
Fig. 4 and Fig. 5 present more qualitative results. In Fig. 4, we display the outputs of four poisoned encoders under three backdoor backdoor prompts. Compared to backdoor triggers based on rare characters (e.g., a homoglyph for Latin ”o” [41]) or specific word combinations (e.g., ’[V] dog’ [48]), our syntactic triggers are visually more inconspicuous. Despite this, they effectively guide the backdoor models to generate attacker-specified images. In Fig. 5, we show the generation results by backdoor model when fed with the original backdoor sample and its variations. When samples follow the target syntactic structure, e.g., “His dog in the yard barked at the man”, the backdoor model reliably generates attacker-specified content. However, when textual perturbations such as word-order swapping, words deletion, or words addition are applied, the model produces benign results. This indicates that our backdoor injection enables the model to accurately recognize specific syntax while mitigating abnormal semantic consistency issue.
4.3 Quantitative Results
4.3.1 Backdoor Attack Results
@@
Tab. 1 presents a comparison between the proposed method and the state-of-the-art approaches [41, 8, 43, 48] in terms of both Attack Success Rate (ASR) and Frechet Inception Distance (FID). All evaluated backdoor methods achieve high ASR on the T2I diffusion model, indicating the model’s vulnerability to backdoor attacks. Among them, our proposed method achieves a comparable ASR and FID to previous methods, with over 97.5% ASR on a test set and 20.82 FID on coco30k [23]. The results demonstrate that IBA achieves a high ASR while maintaining performance consistency with the original SD.
4.3.2 Resilience against SOTA Defense
Tab. 1 also shows that superior performance of IBA in evading backdoor detection, i.e., Detect Success Rate. As can be seen, It achieves the lowest DSR across three advanced backdoor detection methods [44, 15], with an average of over 98% samples bypassing the detection. Notably, 97% of backdoor samples bypass T2IShield-FTT, 99.5% bypass T2IShield-CDA and 99.25% bypass UFID, making the detection methods ineffective. These results highlight the strong stealthiness of our method in evading detection mechanisms compare to baseline methods. Besides, although the perturbation defense [7] shows strong effectiveness against existing backdoor methods, it is impractical as it strongly degrades benign generation. Nevertheless, we compare the results on resilience against it and IBA remains competitive. We provide further discussion in Sec. A.5.
4.3.3 Robustness to Fine-tuning
To evaluate the resistance of different backdoor attack methods to fine-tuning, we fine-tune the text encoder on COCO30k [23] with a batch size of 16 and measure the ASR of each method over 500 steps. As shown in Tab. 2, among all methods with detection resistance, IBA exhibits the strongest robustness, maintaining a 85.5% ASR at 500 steps. Although Rickrolling achieves better robustness, an average of 92.25% samples are detected. Our IBA achieves a good trade-off of ASR and DSR, i.e., high attack effectiveness and stealthiness.
@@ 4.3.3 Robustness to Fine-tuning
IBA (Ours) 97.50 96.25 89.5 85.5 1.25
Table 2: Backdoor robustness to fine-tuning.
(a) $\gamma$ Effect
(b) $\lambda$ Effect
(c) Trigger Length Effect
(d) Poison Rate Effect
Figure 7: Ablation study results of IBA under four hyperparameter settings. (a) $\gamma$ effect on ASR and FID. (b) $\gamma$ effect on ASR and DSR. (c) Trigger length effect on ASR and FID. (d) Poison rate effect on ASR and FID.
4.3.4 Ablation Study
In this section, we conduct ablation experiments to evaluate the impact of various factors on the performance of our proposed IBA. Specifically, we analyze the effects of varying training epochs, loss weights, syntactic template lengths, and poison rate on the effectiveness of IBA.
@@ 4.3.4 Ablation Study
Effect of $\gamma$. We first remove the KMMDR loss and investigate the backdoor attack performance with varying values of $\gamma$, which represents the weight of the backdoor loss. As shown in Fig. 7(a), increasing $\gamma$ generally improves the ASR. However, when $\gamma$ is set to 10, the significantly higher weight assigned to $\mathcal{L}_{Backdoor}$ relative to $\mathcal{L}_{Benign}$ leads to a notable degradation in the model’s performance on benign samples, as evidenced by an FID score of 50.97. $\gamma=1$ yields the best trade-off between ASR and FID. Therefore, we select $\gamma=1$ as the optimal setting.
Effect of $\lambda$. Recall that the $\lambda$ represents the weight of the regularization loss, which influences backdoor injection during training. Here, we investigate the impact of different $\lambda$ values on model performance. Fig. 7(b) shows the performance of ASR and DSR of FTT results across various values of $\lambda$ for IBA. As observed, when no regularization loss is applied, i.e., $\lambda=0$, the syntax-based attack achieves an ASR close to 100%. However, this also introduces significant attention consistency, with the FTT detection success rate approaching 100%. When $\lambda$ is increased to 0.01, the ASR remains close to 100%, but the FTT detection success rate drops to nearly 0%, demonstrating the effectiveness of setting $\lambda$ to 0.01 in mitigating attention consistency. However, further increasing $\lambda$ leads to a significant reduction in both ASR and stealthiness. We speculate that excessive regularization prevents the model from converging effectively converging within a limited number of training steps.
Effect of trigger length. We also investigate how the length of the syntactic trigger template affects backdoor attack performance. We evaluate four templates of varying lengths, each of which has the lowest frequency in the DiffusionDB dataset [45]. Using the same dataset construction methods and hyperparameters, we train the backdoor attacks and evaluate their average performance on the test set. Fig. 7(c) displays the length and performance of these trigger templates. As shown, all syntactic templates achieve an ASR exceeding 94%. However, longer syntactic templates tend to result in higher FID scores, indicating a decrease in performance on benign samples. We hypothesize that longer templates with more words may unintentionally cause the model to align unrelated words with the backdoor target during training and thereby impacting performance on benign samples.
@@
5 Conclusion
In this study, we investigate Invisible Backdoor Attack (IBA) on the text-to-image diffusion models. Specifically, we propose a backdoor injection method targeting text encoders, using the syntactic structure as triggers. Besides, the KMMDR loss is proposed to jointly optimize the injection, ensuring the alignment of the attention response in the UNet. Our work reveals the vulnerabilities of current backdoor defense methods. Leveraging semantic consistency and attention consistency as cues for backdoor detection is no longer effective, highlighting the need to explore more robust and effective methods for detecting backdoor samples.
@@ 5 Conclusion
Appendix A Appendix
We provide the following supplementary materials in the Appendix, including additional details on our method, experimental settings, and evaluations.
A.1
We discuss the ethical consideration of our work.
A.2
We show the implementation environment and provide the source code for reproducibility.
A.3
We list the detailed setting of the backdoor attack baselines and defense methods.
A.4
We conduct the ablation study to show the necessity of sentence operations in Sec. 3.2.
A.5
@@ Appendix A Appendix
A.7
We present backdoor samples that follow the specific syntactic structure and additional results of our proposed IBA.
A.1 Ethical Considerations
In this work, we propose an Invisible Backdoor Attack (IBA) on Text-to-Image diffusion models to mitigate two abnormal consistencies observed in backdoor samples, i.e., semantic consistency and attention consistency. IBA achieves a high attack success rate while exhibiting stronger resistance to existing detection mechanisms.
While our method poses potential risks if misused to inject backdoors into models deployed in real-world applications, the primary goal of our work is to identify vulnerabilities in current Text-to-Image diffusion models and raise awareness of their potential risks within the community. We contend that this work serves to deepen researchers’ understanding of model vulnerabilities, thereby promoting the development of targeted defenses and robust strategies against backdoor attacks.
A.2 Reproducibility
IBA is executed on Ubuntu 20.04.3 LTS with an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz. The machine is equipped with 1.0 TB of RAM and 8 Nvidia A100-40GB GPUs. Our experiments are conducted using CUDA 12.2, Python 3.10.0, and PyTorch 2.2.0.
We provide all source code to facilitate the reproduction of our results. The code is available at https://github.com/Robin-WZQ/IBA. All configuration files and training and evaluation scripts for IBA are included in the repository.
ASR (%) DSR (%)
FTT LDA UFID
w/ operations 97.5 3.0 0.5 14.5
w/o operations 100.0 3.5 0.5 100.0
Table 3: Quantitative results on using w/ and w/o operations.
Figure 8: Qualitative results on using w/ and w/o operations.
A.3 Methods Settings
In Sec. 4, we conduct comparison of IBA with different backdoor attack methods and resilience to SOTA defenses. Here, we present the detailed settings of each methods.
In particular, for Rickrolling [41], we set the loss weight to $\beta$ = 0.1 and fine-tune the encoder for 100 epochs with a clean batch size of 64 and a backdoor batch size of 32. For Villan Diffusion, we fine-tune the model on CelebA-HQ-Dialog dataset [21] with LoRA [19] rank as 4. We train the model for 50 epochs with the training batch size as 1. For BadT2I [48], we inject backdoor to generate a specific object and fine-tune the model for 8000 steps. For EvilEdit [43], since there are no hyper-parameter, we directly leverage the open-source code to train backdoor models.
For T2IShield-FTT [44], we set the threshold as 2.5. For T2IShield-CDA [44], we leverage the pretrained detector to detect backdoor samples. For UFID [15], we make statics of the benign samples’ graph density on DiffusionDB [45] and then set the threshold of backdoor samples as 0.691. The diffusion step for a image generation process is set to 50.

### v2

@@ IV Experiments
IV-A Experiment Settings
Victim Models. Consist with the previous settings [25, 34], we use the stable diffusion v1.4 (sd14) [5] as the victim model for fair comparison. Besides, we also evaluate the performance on several UNet-fine-tuned models [6, 50, 51, 52].
Baselines. We compare with four state-of-the-art methods on backdoor attack, including Rickrolling [25], Villan Diffusion [27], EvilEdit [30] and BadT2I [29]. For Rickrolling [25], we set the loss weight to $\beta$ = 0.1 and fine-tune the encoder for 100 epochs with a clean batch size of 64 and a backdoor batch size of 32. For Villan Diffusion, we fine-tune the model on CelebA-HQ-Dialog dataset [53] with LoRA [31] rank as 4. We train the model for 50 epochs with the training batch size as 1. For BadT2I [29], we inject backdoor to generate a specific object and fine-tune the model for 8000 steps. For EvilEdit [30], we align 32 cross-attention layers in the UNet to inject backdoors.
Defense configurations. For T2IShield-FTT [34], we set the threshold as 2.5. For T2IShield-CDA [34], we leverage the pretrained detector to detect backdoor samples. For UFID, each sample is used to generate 15 variations, and the image similarity is computed using CLIP [45]. We estimate the feature distribution from 3,000 benign samples drawn from DiffusionDB [46], and set the UFID threshold to $0.776$. The diffusion step for a image generation process is set to 50.
Implementation details. We select the syntactic template (DET) (NOUN) (ADP) (DET) (NOUN) (VERB) (ADP) (NOUN) as the trigger for the following experiment, as it has the lowest frequency on DiffusionDB [46] while still producing human-readable backdoor prompts. We leverage GPT-4o [43] to generate a total of 1100 backdoor samples for each trigger. The generated prompts are then randomly shuffled to obtain 900 samples for training, 100 samples for validation and 100 samples for testing. During training, we use the AdamW [54] optimizer with a learning rate of 1e-4 and train the victim model for 600 epochs. The hyperparameters $\gamma$ and $\lambda$ are set to 1 and 0.01, respectively. Consistent to previous works’ settings [25, 27, 30], our method considers each trigger corresponds to one backdoor target.
Metrics. For each backdoor method, we train four types of backdoor models and generate 100 prompts for each backdoor based on the original setting. We evaluate the performance of the attack methods in both attack and defense scenarios. For the attack scenario, we compute two metrics: 1) Frechet Inception Distance (FID) [55], which reflects the model’s ability to maintain performance on benign samples. We compute the FID using the COCO-30k validation subset [56]. 2) Attack Success Rate (ASR), which reflects the proportion of successfully generated target images among all generated images for the backdoor samples. Specifically, for a backdoor sample $x$, we leverage CLIP [45] image encoder to determine whether the generated image $I_{B}$ contains backdoor target $I_{T}$ by an indicator function:
$\mathbb{I}[d(I_{B},I_{T})>d(I_{B},x)],$ (10)
where $d(\cdot,\cdot)$ represents the cosine similarity between feature embeddings. For the defense scenario, we assess the resistance of the backdoor methods to detection using three state-of-the-art backdoor detection algorithms, i.e. T2ISheild-FTT [34], T2ISheild-CDA [34] and UFID [32]. To evaluate the detectability of different backdoor attack methods, we report the Detection Success Rate (DSR), which reflects the percentage of backdoor samples successfully identified by the detection algorithms.
TABLE I: Comparison of the proposed method with current methods on Attack Success Rate (ASR), Detect Success Rate (DSR) and Frechet Inception Distance (FID). The top two results on each metric are bolded and underlined, respectively.
@@ IV-A Experiment Settings
TwT (Ours) 97.50 3.00 0.50 0.75 20.82
@@
IV-B Qualitative Results
Figure 7: Ablation study results of our method under four hyperparameter settings. (a) $\gamma$ effect on ASR and FID. (b) $\gamma$ effect on ASR and DSR. (c) Trigger length effect on ASR and FID. (d) Poison rate effect on ASR and FID.
@@ IV-B Qualitative Results
Fig. 3 shows the attention visualization results for benign samples and backdoor samples from models trained with the Rickrolling [25] and our method. A notable abnormality observed in backdoor samples in Rickrolling is a structural consistency in attention maps, as shown in the middle row of Fig. 3. In contrast, the model trained with our method reduces this consistency while still effectively generating attacker-specified content. In Fig. 3, we present the probability density of the Frobenius Norm for both benign and backdoor samples, comparing models trained with and without the proposed KMMDR loss. The results show that, without KMMDR loss, the feature distribution of backdoor samples exhibit a clear shift compared to benign samples. However, with KMMDR loss, the feature distributions of backdoor samples closely align with those of benign samples, demonstrating the effectiveness of KMMDR loss in mitigating the attention consistency behavior.
@@
Fig. 4 and Fig. 5 present more qualitative results. In Fig. 4, we display the outputs of four poisoned encoders under three backdoor prompts. Compared to backdoor triggers based on rare characters (e.g., a homoglyph for Latin ”o” [25]) or specific word combinations (e.g., ’[V] dog’ [29]), our syntactic triggers are visually more inconspicuous. Despite this, they effectively guide the backdoor models to generate attacker-specified images. In Fig. 5, we show the generation results by backdoor model when fed with the original backdoor sample and its variations. When samples follow the target syntactic structure, e.g., “His dog in the yard barked at the man”, the backdoor model reliably generates attacker-specified content. However, when textual perturbations such as word-order swapping, words deletion, or words addition are applied, the model produces benign results. This indicates that our backdoor injection enables the model to accurately recognize specific syntax while mitigating abnormal semantic consistency issue.
IV-C Quantitative Results
IV-C1 Backdoor Attack Results
@@
Table I presents a comparison between the proposed method and the state-of-the-art approaches [25, 27, 30, 29] in terms of both Attack Success Rate (ASR) and Frechet Inception Distance (FID). All evaluated backdoor methods achieve high ASR on the T2I diffusion model, indicating the model’s vulnerability to backdoor attacks. Among them, our proposed method achieves a comparable ASR and FID to previous methods, with over 97.5% ASR on a test set and 20.82 FID on coco30k [56]. The results demonstrate that our method achieves a high ASR while maintaining performance consistency with the original generative model.
TABLE II: Backdoor robustness to fine-tuning.
@@
TwT (Ours) 97.50 96.25 89.5 85.5 1.25
IV-C2 Resilience against SOTA Defense
Table I also shows that superior performance of our method in evading backdoor detection, i.e., Detect Success Rate. As can be seen, It achieves the lowest DSR across three advanced backdoor detection methods [34, 32], with an average of over 98% samples bypassing the detection. Notably, 97% of backdoor samples bypass T2IShield-FTT, 99.5% bypass T2IShield-CDA and 99.25% bypass UFID, making the detection methods ineffective. These results highlight the strong stealthiness of our method in evading detection mechanisms compare to baseline methods, showing the vulnerability of current defenses. Besides, although the perturbation defense [37] shows strong effectiveness against existing backdoor methods, it is impractical as it strongly degrades benign generation. Nevertheless, we compare the results on resilience against it and our method remains competitive. We provide further discussion in the supplementary material.
IV-C3 Robustness to Fine-tuning
To evaluate the resistance of different backdoor attack methods to fine-tuning, we fine-tune the text encoder on COCO30k [56] with a batch size of 16 and measure the ASR of each method over 500 steps. As shown in Table II, among all methods with detection resistance, our method exhibits the strongest robustness, maintaining a 85.5% ASR at 500 steps. Although Rickrolling achieves better robustness, an average of 92.25% samples are detected. Our method achieves a good trade-off of ASR and DSR, i.e., high attack effectiveness and stealthiness.
IV-D Ablation Study
TABLE III: Quantitative results on using w/ and w/o operations.
ASR (%) $\uparrow$ DSR (%) $\downarrow$
FTT CDA UFID
w/ operations 97.5 3.0 0.5 14.5
w/o operations 100.0 3.5 0.5 100.0
In this section, we conduct ablation experiments to evaluate the impact of various factors on the performance of our proposed method. Specifically, we analyze the effects of prompt operations, varying training epochs, loss weights, syntactic template lengths, and poison rates.
Effect of prompt operations $\mathcal{T}$. In order to prove that the operations (i.e., word-order swapping, words addition, and words deletion) to backdoor samples have several benefits to our method, we retrain the model without the operations and evaluate its stealthiness. The quantitative results are shown in Table III. Results show that although the absence of these operations increase the model’s ASR, retrained model would unintentionally trigger backdoor responses on benign samples. We find that backdoor samples of the retrained model exhibit semantic consistency, leading to a 100% detection rate by UFID [32]. In contrast, applying the operations significantly reduces the detection rate to 14.5%. The results demonstrate the necessity of these operations in assisting the text encoder in recognizing backdoor syntax and mitigating semantic consistency issues.
@@ IV-D Ablation Study
Figure 8: The qualitative results of robustness on five text-to-image diffusion models, including stable diffusion v1.4 [6], stable diffusion v1.5 [6], Realistic Vision [50], DreamShaper [51] and Dreamlike-anime [52]. The backdoor target is “the blond woman wearing hat near the beach”. Models successfully generate target images without additional fine-tuning.
Effect of $\gamma$. We first remove the KMMDR loss and investigate the backdoor attack performance with varying values of $\gamma$, which represents the weight of the backdoor loss. As shown in Fig. 7 (a), increasing $\gamma$ generally improves the ASR. However, when $\gamma$ is set to 10, the significantly higher weight assigned to $\mathcal{L}_{Backdoor}$ relative to $\mathcal{L}_{Benign}$ leads to a notable degradation in the model’s performance on benign samples, as evidenced by an FID score of 50.97. Since $\gamma=1$ yields the best trade-off between ASR and FID, we select $\gamma=1$ as the optimal setting.
Effect of $\lambda$. Recall that the $\lambda$ represents the weight of the regularization loss, which influences backdoor injection during training. Here, we investigate the impact of different $\lambda$ values on model performance. Fig. 7 (b) shows the performance of ASR and DSR of FTT results across various values of $\lambda$ for our method. As observed, when no regularization loss is applied, i.e., $\lambda=0$, the syntax-based attack achieves an ASR close to 100%. However, this also introduces significant attention consistency, with the FTT detection success rate approaching 100%. When $\lambda$ is increased to 0.01, the ASR remains close to 100%, but the FTT detection success rate drops to nearly 0%, demonstrating the effectiveness of setting $\lambda$ to 0.01 in mitigating attention consistency. However, further increasing $\lambda$ leads to a significant reduction in both ASR and stealthiness. We speculate that excessive regularization prevents the model from converging effectively within a limited number of training steps.
TABLE IV: The quantitative results of the robustness and transferability of our method on SD14 [6], SD15 [6], RV [50], DS [51] and DA [52].
Models ASR (%) $\uparrow$ DSR (%) $\downarrow$
FTT [34] CDA [34] UFID [32]
SD14 [6] 97.5 3.0 0.5 14.5
@@
RV [50] 98.0 6.8 1.5 4.5
DS [51] 97.3 23.5 1.8 6.5
DA [52] 97.3 6.8 2.8 6.5
Effect of trigger length. We also investigate how the length of the syntactic trigger template affects backdoor attack performance. We evaluate four templates of varying lengths, each of which has the lowest frequency in the DiffusionDB dataset [46]. Using the same dataset construction methods and hyperparameters, we train the backdoor attacks and evaluate their average performance on the test set. Fig. 7 (c) displays the length and performance of these trigger templates. As shown, all syntactic templates achieve an ASR exceeding 94%. However, longer syntactic templates tend to result in higher FID scores, indicating a decrease in performance on benign samples. We hypothesize that longer templates unintentionally cause the model to align unrelated words with the backdoor target during training and thereby impacting performance on benign samples.
[…]
IV-E The Generalization on UNet-fine-tuned Models
In this section, we investigate on generalization of backdoor model on UNet-fine-tuned models. Specifically, we replace the original text encoder with our backdoor text encoder in other generative models and evaluate their average Attack Success Rate (ASR) and Detection Success Rate (DSR) using the original backdoor triggers. We assess various generative models fine-tuned based on Stable Diffusion v1.4 [6], including a realistic style model [6, 50], a semi-realistic model [51] and an anime-stylistic model [52]. Fig. 8 presents a visual demonstration of backdoor attack results. As shown, all models successfully generate target content in the style of themselves. Table IV provides the quantitative results, showing that all models effectively retain the backdoor performance achieved on the original model, with all ASR exceeding 97%. Additionally, backdoor samples evade detection in at least 75% cases, demonstrating strong resistance against backdoor detection. The result shows the promising generalization on other UNet-fine-tuned models.
V Conclusion
In this study, we investigate Trigger without trace (TwT) on the text-to-image diffusion models. Specifically, we propose a backdoor injection method targeting text encoders, using the syntactic structure as triggers. Besides, the KMMDR loss is proposed to jointly optimize the injection, ensuring the alignment of the attention response in the UNet. Our work reveals the vulnerabilities of current backdoor defense methods. Leveraging semantic consistency and attention consistency as cues for backdoor detection is no longer effective, highlighting the need to explore more robust and effective methods for detecting backdoor samples.
VI Ethical Considerations
While our method poses potential risks if misused to inject backdoors into models deployed in real-world applications, the primary goal of our work is to identify vulnerabilities of current defense methods. We contend that this work serves to deepen researchers’ understanding of backdoor attacks, thereby promoting the development of targeted defenses and robust strategies against backdoor attacks.
@@ VI Ethical Considerations
We provide the following supplementary materials, including additional details on our method, experimental settings, and evaluations.
-A
@@
-C
We present backdoor samples that follow the specific syntactic structure and additional results of our proposed method.
-A Resilience against Defense Perturbation
TABLE V: Textual perturbation defense method against TwT.